1. A block of metal has a width of 3.2 cm, a length of 17.1 cm, and height of 5.7 cm. Its mass is 1.5 kg. Calculate the density of the metal. Express your answer to two significant figures and include the appropriate units (g/cm^3).

a. Calculate the energy of a photon of electromagnetic radiation at 337.1 nm. Express your answer using 3 significant figures.

b. A laser pulse with wavelength 520 nm contains 4.40 mJ of energy. How many photons are in the laser pulse?
Answer #1

1.   Density =

We need to calculate the volume and mass is given =1.5 kg =1.5×1000=1500 g

Also given, Width = 3.2 cm,    Length of 17.1 cm, Height = 5.7 cm

Now we know that Volume = width × length × height

                                Volume = 3.2 cm×17.1 cm ×5.7 cm

                                                    = 311.904 cm 3

Now putting values in density equation we get

               Density =

                Density = 4.80 g/cm 3

a.

Given wavelength (λ) = 337.1 nm = 337.1 × 10-9 m

So energy of photon can be calculated by using formula

E photon =

Where h = Planck’s constant =6.626× 10-34 J×s

C = speed of light =3× 108 m/s

Putting values we get,

E photon =

E photon = 5.89 × 10-19 J

b.

Given, wavelength =520 nm = 520 × 10-9 m

Energy (E) = 4.40 mJ = = 0.0044 J

We know E = nhv

n= number of photons in the pulse

v (frequency) =

So E =

Or n =

Now putting the values we get,

n =

n= 1.15 × 1016 photons